There are 49,000 seats in a sports stadium. Tickets for the seats in the upper level sell for $25, the ones in the middle level cost $30 and the ones in the lower level are $35 each. The number of seats in the middle and bottom levels together equal the number of seats in the upper level. When all of the seats are sold for an event, the total revenue is $1,419,500. How many seats have been sold in the upper level?
Answer by richwmiller(17219)

u+m+b=49000,
25u+30m+35b=1419500,
m+b=u
bottom b = 14400, middle m = 10100, upper u = 24500
We really only need
u+m+b=49000 and m+b=u to solve this for u
2u=4900
u=24500